OpenStudy (anonymous):

5√6x + 8√14x - 9√6x +18 √14x

OpenStudy (australopithecus):

same deal as last time

OpenStudy (australopithecus):

x is just a number

OpenStudy (anonymous):

i got 13√20x +9√8x but it was wrong

OpenStudy (anonymous):

subtract the 5 sqrt 6 to the 9 sqrt 6 add 8 sqrt 14 to 18 sqrt 14

OpenStudy (australopithecus):

\[x(6)^{\frac{1}{2}} + (6)^{\frac{1}{2}} + 5(6x)^{\frac{1}{2}}\] These fractions cannot be added because they are different numbers

OpenStudy (australopithecus):

at least as far as we know, we dont know what number the variable x is

OpenStudy (anonymous):

-4√12x +26√28x ??

OpenStudy (australopithecus):

nvm you are right

OpenStudy (anonymous):

oh haaha lol awesome thnx

OpenStudy (anonymous):

your combining the like terms :)